OpenStudy (anonymous):

Solve: 3(2x - 6) = -4(x - 3) 6 0.6 3 1.8

OpenStudy (anonymous):

You need to distribute so 3 * 2x =? 3 * -6 =? and -4 * x = -4 * -3 =?

OpenStudy (anonymous):

3(2x-6) = -4(x-3) 6x-18 = 12-4x 10x = 30 or x = 3

OpenStudy (anonymous):

You see I was trying to "help" them understand and you just give them the answer. This site is to help people NOT do their homework for them.